OpenStudy (anonymous):

The diagonal of a rectangle is 10. What is the area of the rectangle?

OpenStudy (anonymous):

If it's a "rectangle" then you need more info than that.

OpenStudy (anonymous):

Kk thanx

OpenStudy (anonymous):

50 if the rectangle is a square

OpenStudy (anonymous):

How did you get 50?

OpenStudy (anonymous):

\[10\cos (45)*10\sin (45)\]

OpenStudy (anonymous):

s^2+s^2=10^2 2s^2=100 area of a square is s^2 s^2=100/2=50